OpenStudy (andijo76):

hello my name is emily im using my moms account but i need help on elimination can anyone help me at all?

OpenStudy (anonymous):

what do you need?

OpenStudy (andijo76):

x+y=3

OpenStudy (andijo76):

how would i do that?

OpenStudy (anonymous):

pay attention in school and you wouldnt need help

OpenStudy (andijo76):

um im sorry but i do pay attention -.-'

OpenStudy (anonymous):

um you need two equations to do eliminations, thats what my teacher told me

OpenStudy (anonymous):

if u did u wouldnt need help

OpenStudy (andijo76):

oops x+y=3 x-y=3

OpenStudy (anonymous):

thats better

OpenStudy (anonymous):

you would add the two x's together which gives you what?

OpenStudy (andijo76):

2x+y=3?

OpenStudy (anonymous):

yes. now you you add the three from the first equation to the three from the second equation giving you what? (3+3)

OpenStudy (andijo76):

2x+y=6?

OpenStudy (anonymous):

sorry. i forgot about you

OpenStudy (anonymous):

yes your right

OpenStudy (anonymous):

now you would get 2x=6 because y-y=0

OpenStudy (andijo76):

now dont i divide 2x on each side?

OpenStudy (anonymous):

yes then you would take what x equals and plug it in to both equations

OpenStudy (anonymous):

only plug in your answer to x into x ok? not y

OpenStudy (andijo76):

okay and what i get is what my y is right?

OpenStudy (anonymous):

no. to get your y you would plug in 3(which is what x equals) and subtract it from the three on other side

OpenStudy (anonymous):